OpenStudy (jailbird):

Pre calc Conic Sections, need help. What is the eccentricity of the equation below? (x-5)^2 (y+1)^2 ------ + ------- =1 52 64

OpenStudy (eliesaab):

\[ c^2=a^2 -b^2\\ e=\frac c a \]]

OpenStudy (jailbird):

that first formula is to get the foci points right??

OpenStudy (eliesaab):

You have a and b, find c then e=c/a

OpenStudy (jailbird):

how did you figure that out?

OpenStudy (eliesaab):

sorry a^2=64 and b^2=52

OpenStudy (eliesaab):

\[ \frac{x^2}{a^2}+\frac{y^2}{b^2}=1 \]

OpenStudy (jailbird):

C=√116

OpenStudy (jailbird):

so √116/8 right?

OpenStudy (eliesaab):

\[ c^2=64-52=12 \]

OpenStudy (eliesaab):

\[ c=\sqrt{12} \]

OpenStudy (jailbird):

oops, I added

OpenStudy (eliesaab):

How much is e?

OpenStudy (jailbird):

√12/64 is irrational, but I can match this up to one of the answers, you are awesome.

OpenStudy (eliesaab):

no \[ \frac{\sqrt {12}}{8}=\frac {\sqrt 3}4 \]

OpenStudy (jailbird):

aww, none of my answers match

OpenStudy (jailbird):

oh, over 8, not 8^2, got it

OpenStudy (eliesaab):

What are your answers?

OpenStudy (jailbird):

√3/4 is one of them

OpenStudy (eliesaab):

That is it

OpenStudy (jailbird):

thx